A woman made 7 annual end-of-year purchases of $3500 worth of common stock. The stock paid no dividends. Then after holding the stock for 8 years, she sold all the stock for $38,000. What interest rate did her investment earn?
Answer #1

Answer:

Here, A = 3,500

FV = 38,000

n = 8 years

FV = A(F/A, i, n)

38,000 = 3,500(F/A, i, 8)

(F/A, i, 8) = 38000 / 3,500 = 10.86

From the interest table we find the value of (F/A, i, 8) at i = 8%, is 10.6366 and at i = 9%, it is 11.0285

By interpolation

i = 8% + [(10.6366 - 10.86) / (10.6366 - 11.0285)] * (9% - 8%)

= 0.08 + (-0.2234/ -0.3919) * 0.01

= 0.0857 or 8.57%

Thus, her investment earns an interest of 8.57%.